OpenStudy (anonymous):

simplify (9+4i)(-3-2i)(2-2i) i have nooo idea how to do this!

OpenStudy (anonymous):

Foil two at a time then multiply that answer by last one i^2=-1

OpenStudy (anonymous):

First Outer Inner Last

OpenStudy (anonymous):

take an i out and then mulitply them all together...i think

OpenStudy (anonymous):

(9+4i)(-3-2i) FIRST 9*-3=-27 OUTER 9*-2i=-18i INNER 4i*-3=-12i LAST 4i*-2i=8

OpenStudy (anonymous):

-8i^2=-8*-1=8

OpenStudy (anonymous):

(9 + 4i)(-3 - 2i) = -27 - 18i - 12i - 8i^2 -27 - 18i - 12i + 8 because i^2 = -1 -19 - 30i now (-19 - 30i)(2 - 2i) -38 + 38i - 60i - 60 -98 - 22i

OpenStudy (anonymous):

GREAT!